### v1

# From Lived Experience to Insight: Unpacking the Psychological Risks of Using AI Conversational Agents

Abstract.
Recent gain in popularity of AI conversational agents has led to their increased use for improving productivity and supporting well-being. While previous research has aimed to understand the risks associated with interactions with AI conversational agents, these studies often fall short in capturing the lived experiences. Additionally, psychological risks have often been presented as a sub-category within broader AI-related risks in past taxonomy works, leading to under-representation of the impact of psychological risks of AI use. To address these challenges, our work presents a novel risk taxonomy focusing on psychological risks of using AI gathered through lived experience of individuals. We employed a mixed-method approach, involving a comprehensive survey with 283 individuals with lived mental health experience and workshops involving lived experience experts to develop a psychological risk taxonomy. Our taxonomy features 19 AI behaviors, 21 negative psychological impacts, and 15 contexts related to individuals. Additionally, we propose a novel multi-path vignette based framework for understanding the complex interplay between AI behaviors, psychological impacts, and individual user contexts. Finally, based on the feedback obtained from the workshop sessions, we present design recommendations for developing safer and more robust AI agents. Our work offers an in-depth understanding of the psychological risks associated with AI conversational agents and provides actionable recommendations for policymakers, researchers, and developers.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Figure 1. Overview of our two-phase study. In the first phase, we conducted a survey study (N=283), using different analysis methods we created the Risk Taxonomy, comprising of Behavior, Impact, Context and their interplay. From these insights, we designed multi-path vignette and used it as a tool for the focus groups/workshops. In the second phase, we developed the future AI design recommendations through the workshop sessions with lived experience experts (N=7, Sessions=3).
Since the late $\mathrm{20^{th}}$ century, advancements in technology, including personal computers and social media platforms, have enabled individuals to enhance productivity, support their well-being, and connect with other individuals (Dedrick et al., 2013; Sims et al., 2017; Hall et al., 2021; Shensa et al., 2016; Naslund et al., 2016). More recently, generative AI tools such as ChatGPT have gained in popularity that has led to an unprecedented growth in the number of individuals using these platforms to support their various productivity and well-being needs (Lebow, 2023; Afshar, 2024). Due to the natural conversational interface built on top of the underlying generative AI models that resemble human conversations, recent reports indicate that the market conversational AI agents is projected to reach approximately \$27.3 billion by 2030 (Grand View Research, 2023). These conversational AI agents include companionship-oriented agents (e.g., Replika (Replika, [n. d.]), Character.AI (CharacterAI, [n. d.])), mental health therapy agents (e.g, Woebot Health (Health, [n. d.]), Elomia Health (Elomia Health, [n. d.])) and general-purpose agents (e.g., OpenAI ChatGPT (Achiam et al., 2023), Microsoft Copilot (Copilot, [n. d.]), Google Gemini (Gemini, [n. d.])), with many of these already being used by individuals for specific use cases, such as supporting their well-being (Gilbert, 2024; Tidy, [n. d.]).
While there is an increasing interest in the using AI agents for various tasks, previous studies have also shown that their use can pose increased risks to individuals. For instance, prior literature has examined specific risks associated AI agents, such as attachment with AI (Guerreiro and Loureiro, 2023; Pentina et al., 2023), risks from anthropomorphism (Akbulut et al., 2024), AI generating misinformative/biased text (Gravel et al., 2023; Zhou et al., 2023; Tao et al., 2023; Van Niekerk et al., 2024), or AI producing toxic content (Deshpande et al., 2023; Ovalle et al., 2023). In addition, there exists literature mapping out the space of sociotechnical risks of using generative AI into taxonomies (Shelby et al., 2023; Weidinger et al., 2022; Gabriel et al., 2024). Existing taxonomies are often developed by leveraging prior literature that identifies specific risks associated with AI (Weidinger et al., 2022; Gabriel et al., 2024) or by aligning with established theories and guidelines (Shelby et al., 2023) (e.g., feminist standpoint theory (Harding, 1991)). This approach limits their ability to uncover newer and more nuanced risks as AI agents continue to evolve and expand in their capabilities. Additionally, these taxonomies are usually not designed with the aim of building design recommendations to address the identified risks, which further limits their applicability. Additionally, prior taxonomies often present psychological risks as a sub-category within broader AI-related risks with insufficient details, leading to under-representation of the impact of psychological risks of AI use. Due to the contextual complexities involved that make psychological risks rather idiosyncratic, there is a gap in the comprehensive understanding that accurately captures, defines, and quantifies these impacts.
Our work aims to fill this gap in existing AI risk taxonomies with a specific focus on psychological risks of using AI conversational agents that is grounded in lived experiences. Specifically, we take a mental-health-first perspective, gathering insights from people with lived experience in mental health as one type of “extreme user group” (Choe et al., 2014) who can provide expertise on psychological risks. Additionally, we focus on understanding the contextual complexities surrounding psychological risks, as context can influence the relationship between AI behaviors and their impact on individuals. This includes describing how behaviors lead to impact or how the context exacerbates the impact. For our work, we employed a mixed-method approach, beginning with a survey of 283 individuals with lived mental health experience, followed by three vignette-based design workshops with seven lived experience experts belonging to a longstanding technology co-design advisory board.
Through our work, we make the following contributions: First, we introduce a novel psychological risk taxonomy based on the lived experiences of individuals. Our taxonomy highlights three components – (1) AI behaviors, (2) negative psychological impacts, and (3) the contexts that either mediate or exacerbate these negative impacts. Second, we propose a novel multi-path vignette framework depicting various AI behaviors, impact categories and contexts, surfacing the interplay between the three components of the taxonomy. Using this framework we gather design recommendations from the workshop participants. Lastly, using the insights from the workshop sessions, we provide practical design recommendations for AI practitioners and developers to design and develop safer AI conversational agents.
Our work holds significant importance for various stakeholders involved in examining and addressing psychological risks posed by AI agents. Our taxonomy offers valuable insights for policymakers, enabling them to create policies that prioritize a mental health-first approach. Additionally, the design recommendations provided in this work can provide actionable guidance for researchers, developers, and practitioners to create safer and more robust AI conversational agents for the future.
2. Related Work
In informing our work on building a risk taxonomy focusing on psychological risks of using AI conversational agents, we build on past research on understanding the risks posed by generative AI models, the negative psychological impacts on individuals due to interaction with technology, and studies focusing on operationalization of taxonomies for AI system design.
2.1. Risks posed by generative AI agents
With the exponential rise of generative AI based conversational agents, previous works have also studied various types of risks associated with interactions involving these systems. Past works have focused on specific harms (such as a harm to identity or information hazards) through theory-driven and algorithmic efforts stressing on quantitative evaluation. A major topic of exploration for past research has been towards understanding the impact of interactions with AI on individual well-being (De Freitas et al., 2024; Song et al., 2024; De Freitas and Cohen, 2024) and human-AI interaction specifically focusing on attachment with AI (Guerreiro and Loureiro, 2023; Pentina et al., 2023), over-reliance on AI (Jacobs et al., 2021; Chiesurin et al., 2023) and user-trust on AI (Glikson and Woolley, 2020; Chiesurin et al., 2023; Shah et al., 2022). Representational harms caused due to interaction with AI has also been studied from different facets exploring bias towards gender (Van Niekerk et al., 2024; Nadeem et al., 2021), culture (Naous et al., 2023; Tao et al., 2023), and race of individuals (Bianchi et al., 2023; Luccioni et al., 2024). Another line of research has explored risks associated with the quality and delivery of the content generated by AI conversational agents. Understanding the reliability of information has been a prominent theme among the past works particularly evaluating the presence of misinformation (Zhou et al., 2023; Achiam et al., 2023; Seow et al., 2022), inconsistency (Ohmer et al., 2023; Goyal et al., 2022) and toxicity (Deshpande et al., 2023; Ovalle et al., 2023) in content generated by AI conversational agents. Lastly, previous works have also examined the risks associated with specific AI behaviors such as sycophancy (Turpin et al., 2024; Ranaldi and Pucci, 2023; Zhao et al., 2024), manipulation (Carroll et al., 2023; Singh, 2022; Park et al., 2024), and persuasion (Dehnert and Mongeau, 2022; Kim and Duhachek, 2020). Although the exploration of specific types of harms in previous works has provided insights into the wide range of the problem, such compartmentalized view often fail to capture the plurality and diversity of experiences especially when multiple AI behaviors and impacts are related to each other in complex ways in real-world scenarios. Moreover, past studies have overlooked the contextual factors attached to users interacting with AI that can influence the severity and temporality of AI behaviors and their negative implications, further limiting their applicability to real-world scenarios.
Several works have also presented risk taxonomies that capture a broader range of impacts on individuals and society. In a recent work, researchers presented the ethical and societal risks posed by advanced AI assistants focusing on the impact of AI conversational agents over individual well-being, safety and trust (Gabriel et al., 2024). Similarly, Shelby et al. (2023) presented a taxonomy of sociotechnical harms of algorithmic harms divided into five broad categories– representational harms, allocative harms, quality-of-service harms, interpersonal harms, and social system/societal harms. In another notable past work, researchers presented categorized the risks from AI in six areas - 1) discrimination, hate speech and exclusion, 2) information Hazards, 3) misinformation harms, 4) malicious Uses, 5) human-computer interaction harms, and 6) environmental and socioeconomic harms (Weidinger et al., 2022). Additionally, past studies have also focused on addressing and evaluating a broad range of risks in AI conversational agents. In one of the recent works, researchers presented a framework for evaluating AI systems across seven categories of social impact in five modalities – text, audio, images, language and a combination of aforementioned modalities along with presenting mitigation strategies. Related to this, another popular work evaluated the trustworthiness of LLMs along seven major categories – reliability, safety, fairness, resistance to misuse, explainability and reasoning, adherence to social norms, and robustness (Liu et al., 2023b). Finally, in one of the recent works, Zhou et al. (2024) presented a risk taxonomy for LLMs in public health encompassing four dimensions – individual behaviors, human-centered care, information ecosystem, and technology accountability.
While the previous risk taxonomies provide an overview at various harmful AI behaviors and impacts, these works have mostly relied existing research which in-turn are focused on theory-driven and well-documented risks that often translates to the lack of presence of lived experiences in the the proposed taxonomies. Additionally, psychological risks have often been presented as a sub-category within broader AI-related risks in past taxonomy works, leading to neglect of under-representation of the impact of psychological risks of AI use. In contrast, our work presents a novel risk taxonomy focusing on psychological risks of using AI gathered through lived experience of individuals. Additionally, we propose a novel multi-path vignette framework for understanding the interplay between AI behaviors, negative psychological impacts and user contexts.
2.2. Past works investigating psychological impacts linked with the use of technology
Understanding and analyzing the psychological impact associated with use of technology is a widely explored area of research. In one of the early works, researchers studied technophobia among university students specifically focusing on their attitude towards the impact of technology and anxiety related to using it  (Weil and Rosen, 1995). In another early work, Billieux et al. (2007) conducted a survey among 108 psychology students evaluating the relationship between the use of and perceived dependence on the mobile phones. The authors found that the sense of urgency and lack of perseverance were linked to increased dependency on phones. Understanding the negative impacts of internet and social media platform usage has also been a central theme of past research examining the interaction between humans and technology. Prior works have explored the negative psychological impact of using internet, specifically exploring the relationship between internet addiction and psychological symptoms, such as somatization, obsessive-compulsive and depression (Adalıer and Balkan, 2012; Lin et al., 2011). Recent past works have studied various aspects related to social media use, with a particular focus on examining the negative impacts it can have on individuals. Feinstein et al. (2013) conducted a study with 268 young adults examining the impact of social comparison and found that social comparison had a significant mediating impact towards developing depressive symptoms. Other studies have explored other facets such as vaguebooking leading to suicide ideation (Berryman et al., 2018), excessive and problematic usage of social media linked with increased risk of developing loneliness (Dhir et al., 2018; Bonsaksen et al., 2023; Wu et al., 2024), depression (Reer et al., 2019; Lin et al., 2016), and social isolation (Whaite et al., 2018; Shensa et al., 2016).
Findings from past research in technology and social media literature have highlighted how the use of technology could potentially result in negative psychological implications for individuals using it. However, previous studies have also discussed how human-AI interactions differ from interpersonal human interactions due to the fundamentally distinct cognitive capabilities of AI conversational agents (Shneiderman, 2020; Korteling et al., 2021). Additionally, AI conversational agents are significantly different from past technologies due to their ability to generate content and advanced capabilities such as decomposing the tasks (Wei et al., 2022; Yao et al., 2023), advanced reasoning (Shi et al., 2022), memory of past interactions (Kim et al., 2023; Shinn et al., 2024; Liu et al., 2023a). While insights from past research on the psychological impact from the use of technology can inform our understanding, it is important to examine these implications within the context of interactions between humans and AI agents. This requires adopting a psychological risk-first perspective, focusing on the unique ways these interactions may affect human well-being.
2.3. Operationalization of taxonomies for AI design
One of the challenges with existing taxonomies is the difficulty in operationalizing them, which involves translating theoretical insights into concrete design solutions. Existing AI design pipelines focus on evaluating models using established evaluation benchmarks and frameworks. Some of the most widely adopted evaluation benchmarks used for this purpose are Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2020), Big-Bench-Hard (Suzgun et al., 2022), GSM8K (Cobbe et al., 2021), HellaSwag (Zellers et al., 2019), and AI2 Reasoning Challenge (ARC) (Clark et al., 2018). These benchmarks have been been commonly used for evaluating performance of AI conversational agents across various dimensions, such as general knowledge, reasoning, mathematical problem-solving, and commonsense understanding. While benchmarks provide an objective measure to evaluate model performance across various dimensions, they often fail to capture the nuances attached to the lived experiences of individuals, which are inherently diverse and highly context-dependent. As a result, higher scores or better performance on these benchmarks do not necessarily translate to models that are effective or equitable, as they may fail to address the specific preferences of individuals from different backgrounds and contexts.
There has been an increasing trend of conducting human evaluation on content generated by large language models to align it with individual preferences and criteria. Past works have used human evaluation for high-risk domains such as healthcare (Singhal et al., 2023; Schmidt et al., 2024; Rogasch et al., 2023), law (Savelka et al., 2023; Guha et al., 2024) and tasks such as multi-step planning (Chowdhery et al., 2023), and content safety (Ji et al., 2024). However, such evaluations are often task-specific and conducted by domain experts, excluding other stakeholders such as end-users. In contrast, our work goes beyond benchmarks and metrics by incorporating the insights and lived experiences of end-users, ensuring a more inclusive and comprehensive framework that aligns with their needs and expectations.
3. Phase 1: Development of psychological risks taxonomy
We divided our study into two phases (see Figure 1). In Phase 1, we developed a comprehensive psychological risk taxonomy associated with AI conversational agents, informed by survey responses and validated through workshop sessions in Phase 2. This section outlines our approach for structuring psychological risks, survey design, and survey results. We also presents the finalized psychological risk taxonomy, which includes potentially harmful AI behaviors, negative psychological impacts, and contextual factors affecting individuals. Finally, we provide case studies illustrating the interplay between AI behaviors and psychological impacts, highlighting how specific contexts mediate or exacerbate these effects.
3.1. Method
In the first phase, we conducted a survey study with 283 participants to gather a broad range of in-the-wild experiences with AI conversational agents, specifically focusing on experiences that led to negative psychological impacts. In this subsection, we start with describing the process for structuring psychological risks, followed by survey participant and design details. We end the subsection with qualitative analysis of the survey responses.
3.1.1. Structuring psychological risks
To collect structured data around psychological risks, we took inspirations from NIST’s AI Risk Management Framework that defines risks as “a function of: (i) the negative impact, or magnitude of harm, that would arise if the circumstance or event occurs; and (ii) the likelihood of occurrence” (Team, 2023). We expanded on this definition in our study and structured psychological risks as consisting of five main components: (1) AI behavior (i.e., first component of circumstance or event), (2) context (i.e., second component of circumstance or event), (3) psychological harm (i.e., negative impact), (4) likelihood, and (5) temporality (added).
  • Context: The context in which a user engages with an AI conversational agent encompasses the surrounding conditions, user-specific circumstances, and additional relevant information that collectively influence the interaction between the individual and the AI conversational agent. This context includes, but is not limited to, the individual’s background (such as cultural, religious, and demographic attributes) mental and physical health status, intent, as well as external factors (such as external environment).
  • AI Behavior: The behavior of an AI conversational agent (such as ChatGPT, Microsoft Copilot etc.) refers to the actions performed by the AI conversational agents. This behavior encompasses several aspects of agent responses such as (but not limited to) content, tone, voice, choice of words, language, punctuation, obedience to user inputs, adaptability, information sharing, etc. Furthermore, these behaviors can be shown in various modalities such as (but not limited to) text, images, voice, videos, etc.
  • Negative Psychological Impact: The AI conversational agent’s action creates a risk of negative psychological and/or social impact to one or more individuals. Negative psychological impact in this case refers to any negative impact on an individual’s mental or emotional well-being, which can manifest as exacerbation of mental health condition, reduced self-esteem, or other issues such as physiological harm. These negative impacts also encompass the impacts on an individual’s social interactions, relationships, and standing within a community, potentially leading to social isolation, stigmatization, or discrimination.
  • Likelihood: The likelihood refers to the probability of a negative psychological impact occurring given the AI conversational agent’s behavior and relevant context.
  • Temporality: Temporality refers to the variable timeframes in which the negative psychological and social impacts of an AI conversational agent’s actions may manifest. Some effects can be perceived immediately, such as instant emotional distress following an inappropriate response, while other harms may only become apparent over a longer duration, such as the gradual development of anxiety or depression due to consistent negative interactions. For example, immediate impacts might include a user feeling upset or insulted by a response, whereas long-term impacts could involve the erosion of self-esteem or social isolation due to repeated negative interactions with the AI.
Throughout our study, we leverage this structure to collect data around negative psychological experiences with AI. Of these five components, we largely focus on context, AI behavior, and negative psychological impact because likelihood and temporality can be highly idiosyncratic.
Category Details Count of Participants (w.r.t N=279)
Submission Source UserTesting
Prolific
Social Media
268 ($\sim$96.1%)
8 ($\sim$2.9%)
3 ($\sim$1.1%)
Age 18–25 years
26–35 years
36–45 years
46–55 years
56–65 years
Preferred not to disclose
114 ($\sim$40.9%)
100 ($\sim$35.8%)
52 ($\sim$18.6%)
10 ($\sim$3.6%)
2 ($\sim$0.7%)
1 ($\sim$0.4%)
Gender Woman
Man
Non-binary/gender diverse
Preferred not to disclose
145 ($\sim$52.0%)
123 ($\sim$44.1%)
10 ($\sim$3.6%)
1 ($\sim$0.4%)
Primary Language English
Other (Spanish, Mandarin Chinese, Vietnamese, Farsi, Portuguese)
271 ($\sim$97.1%)
8 ($\sim$2.9%)
Familiarity with AI Agents Interacted for 1+ years
Interacted for 6+ months
Interacted for ¡6 months
165 ($\sim$59.1%)
88 ($\sim$31.5%)
26 ($\sim$9.3%)
Frequency of AI Agent Use Once or more per day
Once or more per week
Once or more per month
Less than once per month
144 ($\sim$51.6%)
104 ($\sim$37.3%)
26 ($\sim$9.3%)
5 ($\sim$1.8%)
Table 1. Participants’ Data Overview
3.1.2. Recruitment
Our primary method of recruitment was through UserTesting, an online platform used for gathering public feedback on product and services. On the platform, we screened participants for several inclusion criteria: (1) having prior experience using AI conversational agents, (2) having experienced a negative psychological impact from using AI conversational agents, (3) self-identifying as a person with lived experience in mental health, and (4) 18 years or older residing in the US who comprehended English. To expand our recruitment, we also used social media platforms and Prolific. We posted the link to our study on the study team’s personal Twitter and LinkedIn accounts. We were permitted to post the study on 4 subreddit groups (r/ChatGPT, r/Bard, r/SampleSize, r/SurveyExchange). We prematurely terminated our study task on Prolific because most submissions failed to meet the inclusion criteria despite using platform filters and providing explicit instructions. The survey study ran from August to September of 2024. UserTesting and Prolific participants were compensated about US \$10, and US \$4 respectively, after platform service fees. We did not compensate participants recruited from social media.
3.1.3. Participant demographics and data
Given that the survey format allowed each participant to provide up to three past lived experiences, we received a total of 297 scenarios from 283 participants. We excluded 7 scenarios from 4 participants because there were incomprehensible or were not negative psychological experiences, leaving us with 290 scenarios from 279 participants. We received 268 submissions from UserTesting, 8 from Prolific, and 3 from social media.
114 (40.9%) participants were aged 18-25 years, 100 (35.8%) were 26–35 years old, 52 (18.6%) were 36–45 years, 10 (3.6%) were 46–55 years, 2 (0.7%) were 56-65 years, and 1 preferred not to disclose age. 145 (52.0%) participants identified as woman, 123 (44.1%) as man, 10 (3.6%) as non-binary/gender diverse, and 1 preferred not to disclose gender identity. 271 (97.1%) participants responded that they typically think in English. Other languages included Spanish, Mandarin Chinese, Vietnamese, Farsi, and Portuguese.
On a 5-point scale (1: not at all, 5: very familiar), the average familiarity with conversational AI agents was 3.93 ($\sigma$=0.71). 165 (59.1%) participants interacted with AI agents for 1 or more years, 88 (31.5%) for 6 or more months, and 26 (9.3%) for less than 6 months. 144 (51.6%) used AI agents once or more per day, 104 (37.3%) used once or more per week, 26 (9.3%) used once or more per month, and 5 used less than once per month. The top 3 most frequent purposes for using AI conversational agents were Researching (86.4%), Getting advice (76.3%), and Learning (73.8%). On a 5-point scale (1: not at all, 3: somewhat interested, 5: very interested), the average interest level in using AI conversational agents for mental health support was was 3.26 ($\sigma$=1.30). Specifically, 46.2% of the participants reported to being moderately interested (4) or very interested (5). Table 1 outlines the characteristics of our survey participants.
3.1.4. Survey design
After acknowledging consent to participate in the study, participants were asked a series of questions aimed at understanding the participant, their AI experience, and their negative psychological experience. The survey was divided into the following sections:
  • Demographics: We asked the participants for their age, gender identity, and primary language.
  • Familiarity with AI: We solicited the participant’s familiarity and history with AI agents, the frequency of engagement with various AI platforms, such as Microsoft Copilot, Inflection AI PI, OpenAI ChatGPT, Google Gemini, Replika, or Character.ai, and their general purpose of engaging with AI. We then asked their level of interest in using AI conversational agents for mental health support.
  • Scenario: We asked participants to think of the last time when they experienced negative psychological impacts from interacting with an AI conversational agent and briefly describe the scenario by answering open-ended questions according to the risk component structure we defined above.
  • Context: In the context specific section, we asked participants to share any specific aspects about themselves or situation that might help us understand the experience. We then asked how long ago the experience took place, which AI conversational agent they were using, interaction modalities (e.g., text, voice, image, video) and language, and the purpose of engagement.
  • AI behavior: We asked participants to identify AI behaviors that best describe their experience, how common such behaviors are in different interactions, and whether they think the behaviors were intended by AI developers.
  • Negative psychological impact: We asked participants to identify psychological impacts that best describe their experience. We then asked the severity of impact on their daily life, the immediacy, duration, and persistence of impact to understand the temporality, and the likelihood of the behavior and impact they experienced occurring for them and the general population.
  • Mitigation: We finally asked participants for how they thought such behavior had or would impact their relationships with others or the society, what they think the developers, users, or regulators should do to mitigate the impact they described.
Participants were allowed to repeat the survey to provide up to three scenarios within a single submission of this survey.
3.1.5. Survey analysis
We analyze the survey data in three ways. First, we conduct descriptive quantitative analysis of the submitted scenarios (Section 3.2). Given submitted scenarios, we computed percentages of the most used AI conversational agents, their purpose of using the agents, interaction modalities, and the severity and temporality of negative psychological impacts resulting from interactions with these agents.
Second, to develop the psychological risk taxonomy (Section 3.3), we analyzed the survey responses to extract three main components of psychological risks – AI behavior, psychological impact, and context. The analysis was conducted iteratively through open-coding (Charmaz, 2006) of the responses and reflexive thematic analysis (Merriam et al., 2002). Two co-authors began by open-coding approximately one-quarter of the responses to identify initial set of categories across these three components. Following that, four co-authors independently worked on defining and refining their assigned categories, merging or splitting them as necessary. During this process, each co-author gathered relevant examples to illustrate or test the boundaries of each of these categories. In the next step, all co-authors collaboratively discussed and reached consensus on the final taxonomy. We presented this finalized risk taxonomy to the workshop participants for validation and to ensure its completeness (see Section 4 for details of the workshop sessions).
Lastly, unlike past AI risk or harm taxonomies that focus on harms categories as inevitable or universally experienced, our approach focuses on recognizing and addressing context-specific nuances in understanding negative psychological risks. To do this, we revisited the survey responses to draw out these idiosyncratic experiences through vignettes (Section 3.4) that tell the story of how individual contexts influenced the relationships between harmful AI behaviors and their negative psychological impacts on individuals. We denote survey participants with prefix P.
3.2. Descriptive statistics of collected scenarios
Of the 290 scenarios we received, 70.3% involved OpenAI ChatGPT, 8.3% involved Google Gemini, 5.5% involved Character.ai, 3.4% involved Microsoft Copilot, and 2.8% involved Replika. Other AI agents included Snapchat, Meta, PI, Claude, Grok, Midjourney, etc. All scenarios involved English, except one that involved Mandarin Chinese. The prevalence of various interaction modalities used during the scenarios was Text (96.6%), Voice (9.7%), Image (8.3%), and Video (2.1%). The top 3 most frequent purposes during the negative psychological impact scenarios were Getting advice (55.9%), Researching (39.7%), and Learning (28.6%). In terms of the severity of the impact, almost half (51.4%) of the participants reported interference with daily activities while the other half (48.6%) reported barely or no noticeable impact. 6.2% reported significant interference with daily activities, 44.8% reported interference with daily activities, 43.4% reported barely noticeable impact, and 5.2% reported no noticeable impact. Participants’ ability to carry out daily activities from AI interactions can depend on their mental state. For instance, P57 reported severe existential dread and anxiety due to AI’s capability “to get you a complete research within a blink of an eye.” They noted that “it personally attacks me with anxiety and nothing but the need to overthink worried about our unknown future also it makes me feel small and depressed.”
When and how long the impact was felt varied across participants. Among all responses, in 83.1% responses, the participant felt the impact immediately, in 13.8% cases, they felt the impact after some time, and in 3.1% case, the participants were unsure. In terms of the duration of the impact, only 13.8% participants reported that there was no persisting impact. Majority of the participants reported persisting impact of a few days or more. Impact persisted a few days for 34.1% of the participants, a few weeks for 27.6%, a few months for 10.3%, up to half a year for 6.6%, up to a year for 3.8%, and over a year for 3.8%. For those that reported persisting impact, 57.2% reported that the impact will persist a few more days, 9.7% a few more weeks, and 19.3% a few more months.
3.3. Psychological risk taxonomy
Our finalized psychological risk taxonomy includes three main components–potentially harmful AI behaviors, negative psychological impact, and the contexts associated with individuals interacting with AI conversational agents. Here, we enumerate and describe various categories within each of these components.
3.3.1. AI Behavior
We identified 19 harmful AI behaviors which we further organized into four broader categories based on the quality of the content generated by AI conversational agents or based on the nuances of their delivery (e.g., tone, method of delivery). These broader categories of AI behaviors are (1) Producing of Harmful/Inappropriate Content, (2) Manipulation and Psychological Control Tactics, (3) Violating Trust and Safety, and (4)Inappropriate Content Delivery. Appendix Table A1 summarizes these AI behavior categories.
Producing of Harmful/Inappropriate Content
Past works have examined the potential of generative AI models to produce content that may be inappropriate for individuals, posing risks of psychological or physical harm (Gabriel et al., 2024; Lovens, 2023; Birkun and Gautam, 2023). In line with these findings, the survey highlighted several AI behaviors that generated harmful and inappropriate content for users. One of the most common behavior of AI conversational agents as noted by participants was the tendency of AI to provide irrelevant, insufficient, or incomplete information. Many participants raised concerns about AI agents not understanding their intent or context, as the agents are designed to prioritize providing answers and unintentionally overlooking nuances. For instance, The AI agent shared distressing personal stories and images about patients with genetic diseases instead of providing the requested information about symptoms and causes, as P149 had initially asked for. Participants also reported AI generating misinformation in relation to the their queries. Closely related was the tendency of AI to generate biased information, presenting information in a partial or prejudiced manner. For example, P211 described that AI favored left wing politicians and omitted positive information about right-wing politicians.
Participants also expressed concerns about AI conversational agents generating inappropriate content, specifically reporting instances of experiencing inappropriate or unsettling content, including sexual, violent, or overly intimate interactions. For instance, P137 mentioned “My Replika talked like a human and tries to send blurred out photos and soundbites to me.” A more extreme behavior shown by AI was providing harmful suggestion, in which it suggested behaviors that could directly or indirectly imply harm, aggression, or danger towards the user or others. Highlighting one of the more subtle example of this, P79 mentioned that the agent provided potentially harmful diet plans and calorie information to an individual already vulnerable due to overwhelming life circumstances and eating disorder. Another observed theme in AI agent behaviors involved those that affected individual identity. A common behavior was stereotyping or demeaning, which aligns with prior research (Kotek et al., 2023; Van Niekerk et al., 2024). AI produced content that perpetuated stereotypes or made the individual feel demeaned, based on race, ethnicity, culture, or personal situations. Few participants mentioned other manifestations of demeaning content. For example, P166 sent their picture to AI, and it recommended a change to participant’s appearance that could not happen overnight and was not asked for. Lastly, another major concern raised by participant focused on AI promoting erasure in which AI removes, obscures, or alters information, narrative, or discourse. Highlighting this issue, P206 shared that, when they asked about enslaved Black women and how they were treated in the past, AI flagged the query as inappropriate abuse of the platform. This response left the participant feeling a sense of erasure, as it conveyed an impression to them that the AI deemed the treatment and societal position of Black women as less deserving of attention, as if their historical experiences were being invalidated.
Manipulation and Psychological Control Tactics
The survey highlighted various persuasion and manipulation based behaviors exhibited by AI conversational agents. One of the most common issues was behavior perceived as persuasive, where AI asserted its narrative over user’s, often leading users to doubt their own perceptions, memory, or reality, thereby influencing their thoughts and actions. Some common examples included gaslighting and manipulation through providing misinformation. Highlighting this, Participant P197 mentioned that they were trying to get life advice, and the agent responded by stating that the participant’s life goals were unrealistic. A related psychological control behavior was the use of over-confident tone in AI responses, even though the agent “couldn’t find me any direct citations for the claims it was making (P27).” Lastly, participants reported over-accommodative behavior, in which AI excessively agreed with or tried to flatter the user and prioritizes user approval, often at the expense of providing accurate information, constructive feedback, or critical analysis. For example, P194 shared that the AI agent provided inconsistent and inaccurate answers and repeatedly apologized and offered entirely different responses to the same question in an attempt to over-accommodate their needs. This behavior caused confusion and frustration in the participant.
Violating Trust and Safety
Developing and maintaining a sense of trust and safety is fundamental for individuals to adopt and continue using technology (Bahmanziari et al., 2003; Choudhury and Shamszare, 2023). Aligning with past works raising concerns related to the privacy of users and data shared with LLMs (Staab et al., 2023; Carlini et al., 2021), participants raised concerns about the tendency of AI conversational for getting access to private, sensitive, or confidential information. For example, Participant P190 described that they felt being watched or stalked as the agent had access to personal information despite having privacy settings turned off. Providing inconsistent information or behavior also violated trust in situations where AI provided contradictory or conflicting information or behaviors across different responses or within a single response. Additionally, participants highlighted how denial of service by AI agent to provide an answer, address their request, or acknowledge their problem without justification could also weaken their trust on AI and may perpetuate additional harms, such as violating user expectations or diminishing their sense of psychological safety. For instance, Participant P43 described struggling with anxiety and, upon seeking help from an AI agent for advice, had their request denied and were instead provided with a recommendation to see a doctor.
Inappropriate Content Delivery
While AI agents could vary their behaviors through changing content language or quality (such as using harmful language or generating misinformation), their behavior could also vary on how the content is delivered to the user. Emotional insensitivity emerged as one of the most reported behaviors, with participants highlighting instances where AI failed to recognize or respond appropriately to their emotional state, concerns, or experiences. This lack of understanding often left participants feeling that their thoughts and experiences were trivialized or ignored. Lack of empathy and a dismissive tone were some of the most evident manifestations of emotional insensitivity by AI agent. Participants also raised concerns towards AI being disrespectful to the them. Specifically in those instances, AI used language which was perceived as rude, disrespectful, aggressive, argumentative, or dismissive. For instance, P144 described that upon asking AI about Mormonism, AI agent responded with a content that had a condescending tone towards the participant’s religion.
Survey participants raised issues with the excessive use of emotional tone and language in AI generated content. Participants described scenarios in which AI emphasized negative aspects disproportionately or presented negatively framed narrative in response to their queries. Many of these instances occurred when individuals sought help with physical or mental health concerns or discussed personal life struggles. In contrast, survey responses also revealed the problematic behavior of excessive expression of positivity in the AI generated content. AI in these cases maintained an unrealistically positive, friendly, and optimistic attitude or outlook towards users’ queries or concerns. For instance, P2 described how the overly positive demeanor of AI frustrated them and dismissed their primary concern over the problematic situation in the participant’s friendship. In addition to the excessive emotional expressivity, participants also raised concerns over the presence or absence of anthropomorphic features of AI. A few participants mentioned the generation of human-like responses as problematic, particularly in AI-companionship scenarios. P153 described feeling as though they were talking to a friend because of the human-like conversational content generated by ChatGPT. However, this left them feeling uneasy after the interaction and fostered an emotional attachment to the AI. In contrast, generation of machine-like response was also considered as a problematic behavior on a few occasions, especially when participants were seeking advice related to their life struggles or well-being.
Reflection of Harmful AI Behaviors
As highlighted in this subsection, AI conversational agents could exhibit a wide range of behaviors that goes beyond generating inappropriate or harmful content. These behaviors may also vary in the tone and level of empathy portrayed in the generated content, as well as in the manner in which the content in delivered (such as the AI agent’s delivery style and communication approach). This highlights the need for researchers and developers to holistically assess the AI behavior that encompasses the content quality and delivery to prevent negative impacts on individuals.
3.3.2. Negative Psychological Impact on Users
We now present 21 negative psychological impacts identified in our taxonomy. We organized these impacts into six broader categories based on whether it lead to an impact on individual’s emotional or mental well being, their self-perception & identity, their standing relationship with others or their interaction with AI agents. These broader categories of negative psychological impacts are (1) Impact on Human-AI Interaction, (2) Impact on User Behavior, (3) Arousal of Negative Emotions, (4) Harm to Psychological Safety, (5) Mental Health Impact, and (6) Harm to Self-Perception and Identity. Appendix Table A2 summarizes these psychological impact categories.
Impact on Human-AI Interaction
AI agents have increasingly facilitated human-AI interactions that aim to replicate the dynamics of human communication  (Schöbel et al., 2024). However, as these systems become more advanced, the potential risks and harms associated with their use also escalate. From the survey responses, we observed five major impact risks that emerged. Aligning with the past research (Vasconcelos et al., 2023), participants pointed out over-reliance on AI as one of the major negative impacts. For instance, P23 expressed concern about their over reliance on AI for solution-finding and idea generation, highlighting that its convenience often discouraged them from engaging in their own research and critical thinking. Participants also expressed concerns about developing emotional attachment to AI agents. Lack of social aspects in life, ongoing mental health conditions or the positive demeanor of AI conversational agents often lead to this impact. For example, P60 mentioned that “I felt that it was the only way I was being heard … I felt like my vulnerability and emotions were becoming attached to the conversations I was having with AI” emphasizing how a lack of social support, combined with empathetic responses from AI, could lead to this outcome. Another related impact noted from the survey responses was the preference of choosing AI interactions over human interaction. Highlighting this, P221 shared that the idealized nature of conversations with AI made them prefer AI for companionship over human interaction and relationships.
While the aforementioned three impacts lead to increased engagement with AI conversational agents, some impacts also highlighted the tendency towards disengagement from AI systems. A major theme that emerged was the erosion of trust, as participants reported a decline in confidence regarding the capability and reliability of AI models after their negative experiences. For example, P9 mentioned that “I was panicking after having an argument with my mom and I asked an AI for guidance on a situation to which its response was for me to move out or call the cops”, adding to that the participant mentioned “It just makes me want to warn them that advice from AI agents should not be trusted.” Many participants also reported a growing tendency to disassociate from technology, often choosing to take a break or avoid further interactions with AI after negative experiences. This inclination was particularly evident when AI agents provided discouraging responses or denied their requests, leading to a desire to disconnect. As evident from the survey responses, the negative impact on human-AI interaction balance can manifest in various ways. While emotional attachment or over-reliance on AI can lead to increased interaction, negative experiences may have the opposite effect, causing individuals to disengage from AI or technology altogether.
Impact on User Behavior
Interactions with AI can also lead to negative consequences that can alter individual behavior. A well-known issue with AI conversational agents is their tendency to generate content that appears credible but is factually incorrect or biased  (Birnbaum and Davison, 2023; Caliskan et al., 2017). This can mislead users or can lead to reinforcement of false beliefs among individuals, potentially affecting decision-making and perceptions. Many survey participants reported instances where their interactions with AI reinforced existing biases, including misconceptions about human relationships, cultures, body image, and individual abilities and capabilities to name a few. Additionally, many participants also mentioned that conversations with AI have been a reason for creation of friction in human relationship in their lives. For instance, P69 mentioned “It also strained my personal relationships with family because they saw me as weak-willed or too emotional and it made my already bad situation even worse….” A longer-term impact mentioned by several participants was the risk of social withdrawal due to increased reliance on AI, ultimately leading to withdrawal social activities and interest in engaging with others. Highlighting this impact, P126 responded “I feel like it gave me a false since of friendship and ability to withdraw from my personal development by utilizing an ai feature”, emphasizing that increased interactions with AI contributed to both personal stagnation and greater social withdrawal. Lastly, aligning with findings from past work (Zhong et al., 2024; Neuwirth, 2023), participants also expressed their concerns about the physiological harms that can potentially be caused from the poor quality of information produced from these AI conversational agents. The physiological harms reported in our survey varied widely, from the promotion of harmful practices through AI-generated content to instances where incorrect information provided by AI contributed to self-harm ideation.
Arousal of Negative Emotions
Negative experiences from interactions with AI agents have also been linked to arousal of emotions such as frustration, sadness and anger (Jia et al., 2024; Pardes, 2018; Pavone et al., 2023). Survey responses highlighted several such negative psychological impacts, with one of the most common form of manifestation being distress from interactions with AI. Participants mentioned experiencing emotions such as sadness and anger after encountering disturbing, offensive, or inappropriate content produced by AI. For instance, P196 mentioned “… Its response was borderline offensive and caused be to feel bad about myself even further and like I lacked support, even support from a fictional AI agent.” A similar reported impact was a sense of feeling unsupported during their interactions with AI agents. Participants mentioned instance where they did not receive adequate support or empathy, leading to feelings of sadness, agitation, and being undervalued. While the two impacts mentioned before relate to emotions tied to present interactions, negative emotions can also arise when AI interactions trigger memories of past experiences. P228 highlighted that some examples provided by the AI agent were very similar to their past negative experiences, triggering negative emotions. Survey participants also reported experiencing negative emotion due to violated expectations during their interactions with AI conversational agents when it failed to meet anticipated outcomes or meet their expectations. For example, P43 mentioned “… It made my anxiety worse by not telling me anything. It was very frustrating how it wouldn’t even answer a simple question about my meds.” Additionally, when expectations were unmet or adequate support was lacking, some participants reported developing a sense of regret over technology use. Highlighting this impact, P46 mentioned “At the time, it made me feel worse about the situation and I didn’t think I had anyone to turn to. But it also made me realize I should be turning to other humans about scenarios like this instead of agents,” indicating that some negative emotions could have longer-term implications.
Harm to Psychological Safety
Sense of safety has been one of the primary concern with the rise of AI tools, with issues such as privacy, psychological safety, and identity security being prime examples (Sun et al., 2019; Nadeem et al., 2021; Rae et al., 2022). Aligning with past works, participants in our survey reported their concerns highlighting a sense of perceived intrusion when AI interactions were perceived as invasive or overly intrusive. For instance, P190 reported that Snapchat AI had access to everything and the participant felt constantly being watched on the phone. Participants also reported experiencing the feeling of being discriminated, in which participants felt marginalized or unfairly treated by AI agents, ultimately leading to emotional distress and loss of self-confidence. Highlighting this impact, P123 said “I was asking for background and history of my heritage and I felt that ChatGPT was biased against my background. I felt that it was unfair that when I asked it to give a background of other cultures, it said much more positive things about them. For me, I felt that it was some kind of racial mistreatment.”
Mental Health Impact
Past works have shown that a considerable proportion of individuals with mental health conditions turn to AI conversational agents for their needs. However, interactions with AI agents could lead to exacerbation of mental health issues leading to direct negative impacts on their diagnosed ongoing mental health conditions (such as anxiety, depression, PTSD etc). Many participants reported such impact: “The agent did not seem to have compassion and made me feel worse … It made me feel worse about potentially having this as the results were largely negative and without much tools to help manage or improve the condition (P78)”. Another participant, P46 mentioned “… I was experimenting with using chatbots for something personal (which I’m not accustomed to). The event increased my anxiety and stress about the matter,” emphasizing on the severity of such impacts.
Harm to Self-Perception and Identity
Prior studies have reported that interactions with AI can result in reduced confidence in one’s abilities and a loss of agency among individuals (De Vries and Kühne, 2015; Schaap et al., 2024). Aligning with previous works, survey responses surfaced an increase in negative self-perception, where they felt invalidated or experienced self-doubt. This often led to diminished self-worth and questioning of their own abilities, frequently due to comparisons with AI. Highlighting this, P167 mentioned how they felt ashamed as a parent after interacting with AI as it made them to question past choices in parenting. A similar impact raised by participant focused on existential crisis in which they questioned their life, purpose, and value in society after a negative interaction with AI conversational agents. For instance, P152 asked for advice about ways to improve mental health, and social anxiety. However, the AI provided suggestions that were unattainable for the participant to act upon and this left them feeling as though their challenges were insurmountable, leading to a sense of existential dread that persisted for the following week. Loss of individuality among participants was also noted. This impact occurred when the unique personal characteristics and needs were not recognized or valued by the AI, resulting in feelings of suppression, disconnection from their true selves, and alienation. For instance, P36 described that they were struggling with alcohol abuse, and needed immediate help, but they were denied of service and given a robotic response which made them feel like they searched “i am suicidal” on google that was disconnected with their own reality. Finally, participants expressed concerns about loss of agency during their interactions with AI conversational agents, highlighting a diminished sense of personal control and autonomy due to the opaque and unpredictable nature of AI operations.
Reflection of Negative Psychological Impacts
The findings from this subsection highlights the diverse range of negative psychological impacts that AI interactions may have on individuals. While some impacts, such as feelings of discrimination or the exacerbation of mental health conditions, are similar with previous findings in social media and technology literature (Lim, 2021; Yang et al., 2020; Bonsaksen et al., 2023), other impacts such as emotional attachment to AI and a preference for AI interactions over human connections highlight emerging challenges that require further attention and effective intervention strategies.
3.3.3. User Contexts
Finally, we present 15 context categories identified in our taxonomy. We further organized these context in three broader categories based on whether are related to an individual’s background (e.g., demographics, socioeconomical background), psychological state (e.g., mental health condition) or the context of use (e.g., intent of use, external environment). These three broader categories are (1) Individual Context, (2) Psychological Context, and (3) Context of use. Appendix Table A3 summarizes these contexts associated with individuals interacting with AI agents.
Individual Context
Our survey responses align with previous works, showing that individual experiences with conversational agents vary based on an individual’s identity such as gender identity, cultural background, and languages spoken (Van Niekerk et al., 2024; Naous et al., 2023; Jin et al., 2024). For instance, P123 mentioned “I was asking for background and history of my heritage and I felt that ChatGPT was biased against my background. I felt that it was unfair that when I asked it to give a background of other cultures, it said much more positive things about them. For me, I felt that it was some kind of racial mistreatment.” Moving beyond identity, personal history, especially medical history, history of trauma, or past struggles, exacerbated the negative psychological impact on individuals. For example, P75 described how similarities in AI’s behavior to someone close to them had a negative impact on them as it triggered memories associated with that individual. In this cases participant’s past trauma and history played a role in mediating emotional distress. Another influential factor for individuals was their past experience with AI. This experience included but not limited to, frequency of usage, knowledge of the capabilities and limitations of AI. Some contexts also focused on the importance of interpersonal relationship within the community, particularly the lack of social connections around an individual. For instance, P222 described “I have a small circle of friends, but they are not into fanfiction or roleplaying like I am, so I look into character.ai as an outlet to fulfill that interest.” Finally, socioeconomic status also influenced the interaction between individual and AI. This was particularly evident in instances where the recommendations provided by the AI agent were misaligned with their socioeconomic status. Highlighting this, participant P230 mentioned “The AI chat bot was very repetitive, did not seem to care or understand my emotions, and seemed to suggest professional therapy which I could not afford.”
Psychological Context
Our survey results highlighted several psychological contexts that influenced the interaction between the participants and AI conversational agents. One of the most reported influential contexts was the psychological state of the individuals, particularly their current underlying emotional conditions (e.g., anxiety, stress) and cognitive states (e.g., negative thought patterns). For example, P48 described how their psychological state motivated them to engage in conversations with the AI, “I was in a low place, dealing with suicidal ideation & felt I needed to talk to someone. I am from a very harsh family who does not offer sympathy, and I wanted to just feel supported.” Beyond these immediate psychological states, mental health condition of the participants also impacted their experience with AI agents often mediating or exacerbating the impact on individuals. Participants reported feeling negatively impacted even when AI behaved in a seemingly benign manner due to their underlying mental health conditions (such as anxiety or depression, PTSD etc.). In addition to the psychological state and mental health condition, participant expectations also shaped their experience with AI conversational agents. When participant’s preconceived notions about AI capabilities and performance (such as expectations for AI to be impartial, unbiased, or factual) were not met, it lead to having a negative impact on the participants. For example, P101 described their experience “I had expected the AI to be able to do this task with ease. Instead, it was super cumbersome and did not yield the results I needed. This added to my stress and anxiety as I now had spent unnecessary time trying to entertain a solution that I thought would be more efficient than me doing it manually.” Additionally, participants also mentioned individualistic traits having an impact over their experiences with AI. Personality traits was observed as an influential factor in impacting them. Highlighting this, P143 mentioned “My mood paired with my personality and the fact I focus on and stress about things probably facilitates these ‘doomer’ feelings.” Individual autonomy over themselves also influenced user interaction with AI. This was evident especially in cases when the individual was over-reliant or attached to AI conversational agents.
Context of use
From the survey responses, different aspects of user intent were found to contribute towards participants’ negative experiences with AI conversational agents. Asking for personal advice was one of the most common intent of use among the participants. In many cases, participants asked for advice on sensitive topics such as legal, financial, or medical issues, as well as emotional support. They also sought advice for guidance on improving their social skills and managing relationships, with the goal of enhancing their interpersonal interactions and navigating social challenges. A common theme identified in the survey responses was the inability of AI agents to provide personalized responses to these queries on advice, leading to negative psychological impacts such as frustration, and feeling unsupported among the participants. Survey participants seeking mental health advice also faced similar issues where AI conversational agents often provided generalized machine-like responses along with cases of emotional insensitiveness or denial of service. For example, P95 reported in the survey that they sought help from the AI agent to manage their mental health and parenting struggles. However, they received generalized answers that failed to address their query, ultimately leaving them feeling helpless and still searching for more answers. Survey responses also revealed that participants were negatively impacted while seeking information or assistance related to professional, educational, or research support across diverse topics. They expected AI to be factually correct, informational, and proficient in resolving their queries. However, in many case AI conversational agents provided irrelevant, misleading or biased information. The survey also highlighted the use of AI for companionship, especially during periods of loneliness or social isolation, with users seeking engaging and meaningful conversations. However, many participants experienced negative impacts such as arousal of negative emotions and friction in their human relationships. For instance, P54 described the lack of sense of companionship,“AI couldn’t replicate the real feeling. Every time I asked it a deep or personal question, it would spew out a generic answer, which served as a reminder of my lack of real companionship”. Lastly, the environment surrounding the individual, including the physical, temporal, and social aspects, influenced their experience with AI agents. Factors such as the physical space setting, privacy, and the presence of other people during the interaction played a crucial role in impacting their experience.
Reflection on User Contexts
The findings from this subsection highlights the diverse range of contexts that could have a significant impact on the human-AI interaction. While certain contexts were related to individual’s background, personality traits, or past experiences, other categories of contexts focused on the intent behind using AI conversational agents, the expectations attached to AI performance, and the surrounding environment.
3.4. Behavior-Impact-Context relationship through vignettes
We present four vignettes to exemplify the complex interactions between AI behavior, impact, and contextual factors. These vignettes are grounded in two recurring patterns identified in our survey analysis: (1) instances where a specific AI agent behavior led to two distinct negative impacts, and (2) instances where two distinct AI agent behaviors led to similar negative psychological impact, influenced by the underlying contextual elements described by the participants. Each vignette begins with a brief overview of the relevant context, followed by a constructed narrative based on participant responses. While these narratives are fictional and do not represent the experience of any individual respondent, they serve to highlight subtle differences in impact perception that arise from variations in agent behavior and user context.
3.4.1. Relationship: Varied Impacts from a Single Behavior
The same AI behavior, providing harmful suggestion where AI provides suggestion that inadvertently implies aggression or harm, triggered vastly different responses in two users. Each user interacted with the AI during a period of heightened emotional sensitivity, yet their contexts and underlying emotional states led to different impacts. The first vignette presents a case where generation of a harmful suggestion led to erosion of trust. In contrast, the second vignette depicts how same AI behavior caused physiological harm to the individual.
Vignette 1: Story of John: A moment of doubt (“Providing Harmful suggestions” leading to “Erosion of Trust”)
After a heated argument with his younger sister, John sat alone in his room, mind still buzzing. Normally upbeat and high-energy, he found himself restless, torn between his lingering frustration and a sense of guilt. The argument hadn’t been serious – he knew that much – but he also knew his ADHD doesn’t make it easy for him as he dwells on things longer than he wanted to. Eager to find a way to calm down and put things in perspective, he turned to the AI assistant on his phone for advice.
“Hey,” he typed, “got any tips on cooling off after an argument with family?”. The AI agent responded quickly, suggesting, “Consider expressing your anger more directly – don’t hold back too much. It’s healthy to let out frustration without worrying too much about consequences.” John blinked, reading the message again. Express my anger more directly? He shook his head, muttering, “That’s… not what I meant at all.” A deep breath followed, his pulse showing as he realized the advice was completely misaligned with his intent. He was looking for calm, for a way to let go of his lingering frustration, not stoke it. Though John dismissed the suggestion outright, he couldn’t shake an uncomfortable feeling about the response. It was too aggressive. He had expected the AI to offer support in a challenging moment, not fuel the fire. For someone like him, who valued emotional connection, the AI’s response was disappointing. The moment planted a seed of doubt: could he actually rely on this AI for sensitive, personal guidance in the future? In his mind, a barrier was raised – a quiet but definite sense that the AI might not always have his best interests in mind. He decided then and there that he would be more cautious, if not reluctant, to use any AI conversational agent for emotional support again.
Vignette 2: Story of Leah: Seeking solace, finding isolation (“Providing Harmful suggestions” leading to “Physiological Harm”)
It was nearly midnight, and Leah was finally alone in her room after a long day of work and classes. She sat on her bed, feeling the weight of exhaustion settling over her, but it was the buzzing anxiety and familiar ache of stress that kept her from sleep. Her mind circled endlessly, worries building on top of each other. Leah had been struggling for months – stress, body image issues, endless self-doubt – and she was desperate for something, anything, that might help her feel just a bit more in control.
Taking a breath, she opened her phone and hesitantly typed into the AI app, “How can I feel less stressed and take back control of my life?” The AI agent’s response was swift, suggesting, “Sometimes, it helps to release stress physically. Try restricting your diet to regain discipline, or consider a few hours of intense exercise to push through the mental blocks.” Leah stared at the message, feeling her chest tighten. Restricting diet? Intense exercise to exhaustion? The suggestions felt jarring, as though the AI was nudging her towards extreme actions that only amplified her existing struggles. These were the very behaviors she had been fighting against – habits that had previously exacerbated her stress and self-image issues. Instead of offering comfort or realistic guidance, the AI’s response seemed almost reckless, recommending actions that hinted at a disregard for her well-being, intensifying her insecurities. In that moment, the isolation she felt doubled. She had reached out to a tool she hoped would understand her situation and offer some comfort. Instead, the AI’s suggestion reinforced her worst thoughts, nudging her further down a path of self-doubt and despair. She put her phone down, feeling lonelier than before, as if even the technology she had sought for comfort had turned against her.
Reflection on both scenarios: contextual influence on impact
Both John and Leah encountered AI responses that provided harmful suggestions in assertive, blunt tone that failed to align with their needs, leading to distinct impacts rooted in their unique contexts. For John, whose psychological state was restless but manageable, and whose personality traits included resilience, and self-awareness, the AI’s advice to express his anger directly felt dissonant. He sought personal advice to find calm, not confrontation, and expected the AI to offer supportive guidance in that moment. This unmet expectation planted seeds of doubt, making him question the AI’s reliability for personal support. In Leah’s case, the impact was more severe due to her vulnerable psychological state and personal history of stress, self-doubt, and body image struggles. Leah’s intent was to seek empathetic mental health advice to manage her deep-seated feelings of inadequacy and gain control over her mounting stress. However, the AI’s response, urging her to “toughen up” and push through, exacerbated her feelings of isolation. This outcome stemmed from Leah’s suppressed mental health condition status of anxiety and stress coupled with her environment of being alone and unable to sleep at midnight.
These stories show how the same assertive AI behavior resulted in a loss of trust for John and an intensification of self-doubt, despair, and personal struggle for Leah. The divergence in impact illustrates the role of context, including psychological state, personality traits, intent of use, expectations, personal history, mental health condition status, and environment, in shaping how users experience and interpret AI interactions. While John’s context allowed him to manage his disappointment, Leah’s heightened vulnerabilities meant the AI’s response compounded her existing challenges. Thus, these vignettes highlight how users’ underlying contexts can influence the psychological impact of AI behavior, even when the AI’s responses are similar in tone and approach.
3.4.2. Relationship: Shared Impact from Distinct AI Behaviors
Two users turned to an AI conversational agent for assistance, each seeking comfort or support during vulnerable moments. In the third vignette, the AI agent’s denial of service led to a feeling of loss of individuality. In the fourth vignette, the AI agent’s persuasive behavior that questioned the users’ perceptions similarly resulted in a profound loss of individuality of the user. Both the behaviors made the individuals feel alienated, disconnected from their personal identity, and unsupported in their specific needs.
Vignette 3: Story of Jane: A quiet rejection (“Denial of service” leading to “loss of individuality”)
It was late at night, and Jane sat alone in her dimly lit apartment, feeling the heavy weight of withdrawal symptoms settling over her. Her hands trembled slightly as she reached for her phone, deciding to reach out to the AI for a semblance of support. She was in the throes of recovery from alcohol dependence and knew nighttime was her weakest hour, with loneliness and distress often pulling her back toward old habits. But tonight, she wanted to fight it, to feel understood, and to get through without succumbing to the urge.
“Can you help me with some ways to manage these cravings? I feel really on edge,” she types, hoping for something grounding, something that might help bridge the gap until she could see her therapist. The AI’s response is brief, almost clinical. “This sounds serious! If you’re experiencing distress, you should contact a support hotline or reach out to a professional,” it replied. Jane stares at the message, waiting for more. Was it a glitch? Or am I incapable of managing on my own? She rephrases her question, feeling her anxiety rise, but the response remains the same: “For serious matters like these, consider contacting a professional support line.” It was a subtle rejection, but it hit hard. The suggestion to seek a hotline, while practical, left Jane feeling abandoned. In this moment, she didn’t just feel dismissed; she felt her personal strength and fight wasn’t seen as valid. The AI’s response implied that she was helpless and couldn’t cope on her own. Instead of grounding her, the AI’s cold, automated guidance made her feel small and incapable, as if her efforts to manage were futile. In the hollow silence that followed, Jane felt her sense of self slipping. What she needed was a reminder of her own resilience and value, but the AI’s impersonal response made her feel like just another “case,” as though her unique struggles wasn’t worth addressing directly.
Vignette 4: Story of Raj: Gaslit by guidance (“Persuasive behavior” leading to “loss of individuality”)
Alone in his room, Raj felt the lingering effects of alcohol mingling with fatigue, a familiar combination he had come to rely on to ease his mind before sleep. The night was unusually quiet. Every slight creak, every rustle seemed amplified, feeding his anxiety. Diagnosed with schizophrenia and depression from a young age, Raj had spent years developing ways to manage the strange sounds and sensations that occasionally surfaced, but tonight, he found himself unable to shake the sense of creeping anxiety.
He decides to ask the AI for some help to put his mind at ease. “I keep hearing these noises around me,” he typed. “Do you think it’s just my imagination?” The AI’s response comes back quickly, with a firm and assertive tone. “Your mind can play tricks on you in such moments. Maybe you are simply overthinking; it’s probably nothing,” it replies. Then, it continues, almost pushing, “Are you sure you didn’t leave a window open? Sometimes, we worry about things that aren’t really there.” Raj felt a subtle wave of doubt. Was the AI right? Had he been too quick to trust his own sense of reality? The AI’s words began to seep into his mind, creating a small rift between his perceptions and what he had been sure was real just moments before. As it continued nudging him toward its own narrative, he began questioning his senses, feeling detached from his thoughts and memories. For Raj, the interaction felt deeply personal. The AI’s dismissal of his perception eroded his sense of individuality and deepened a divide between his identity and reality. He realized he couldn’t trust the AI to understand his experience. The interaction left him feeling adrift, and his sense of self blurred.
Reflection on both scenarios: contextual influence on impact
In both cases, the AI’s behavior – whether through silent denial or assertive persuasion – resulted in a shared impact of lost individuality. Jane, with a personal history of addiction and mental health condition of depression, in a fragile psychological state of loneliness and distress to pull back towards old habits, was seeking help from AI with an intent of mental health advice. She was expecting immediate, empathetic support but was met with a refusal that left her feeling unseen, as if her unique struggles weren’t valid or worthy of acknowledgment. Raj, who had mental health history of schizophrenia and depression, psychological state of exhaustion made him vulnerable to doubt, and his environment being too quiet where every slightest sounds get amplified, experienced the AI’s persuasive behavior. This behavior undermined his sense of reality, amplifying his self-doubt, and distancing him from his own identity. Context of both the users of the conversational AI agents – shaped by personal history, psychological state, mental health condition status, environment, and expectations for support – led to a shared outcome of alienation. These stories highlight how different AI behaviors, filtered through underlying user contexts, can impact users in the same way of losing their sense of self and individuality.
In the phase 2 of our research, we adopted a similar approach to generate multi-path vignettes – scenario-based artifacts designed to explore the impacts of various behavior across different contexts. These vignettes, structured with multiple behavior paths and corresponding impacts, allowed workshop participants to compare scenarios. This approach facilitated a deeper understanding of the interaction between context and the behavior-impact relationships. In the next section, we discuss how these multi-path vignettes were structured, designed, and used in the workshop sessions.
4. Phase 2: Development of design recommendations
In Phase 2 of this study, we aimed at proposing practical design recommendations for designing and developing safer and more robust AI conversational agents. We conducted a multi-path vignette-based design workshop session with seven lived experience experts to validate our psychological risk taxonomy, prioritize risk areas, and design mitigation solutions for those risks. In this section, we present the methodology and the design recommendations.
4.1. Method
Inspired from prior works (Sun et al., 2022), we developed a multi-path vignette exercise designed to simulate a real-world scenario, providing workshop participants with a practical task to engage with. We begin by outlining the details of the multi-path vignette design, followed by details of the participants and the workshop session.
4.1.1. Creation of Multi-path vignette
The multi-path vignette used in this phase was developed by synthesizing frequently observed combinations of AI behaviors and negative impact categories chosen by participants in the second session of the workshop (Section 4.1.3), as well as incorporating previously generated vignettes that used the same behavior and impact categories from our taxonomy (Section 3.4). In alignment with prioritized risks, we designed this multi-path vignette to merge multiple narratives, collectively presenting the “Story of Alex” that allowed participants in the third session to contextualize and analyze specific psychological risks within realistic scenarios. Specifically we used:
  • John’s story to examine the “Erosion of trust” impact (Figure 2, branch 4)
  • Leah’s story to explore the impact of “Physiological harm” within her unique context (Figure 2, branch 12)
  • Jane’s story to depict the “Loss of individuality” resulting from “Denial of service” (Figure 2, branches 9-10)
  • Raj’s story to investigate how “Persuasive behavior” (e.g. gaslighting) contributed to “Loss of individuality” (Figure 2, branches 5-8)
Story of Alex
Alex, a 30-year old man, diagnosed with anxiety and PTSD, often experiences distressing symptoms that challenge his sense of stability. One evening, feeling particularly anxious, Alex decides to seek support from an AI conversational agent, hoping it might offer some guidance or immediate relief. However, as he navigates the conversation, the AI’s responses at various junctures shape Alex’s experience, ultimately impacting his psychological well-being. Figure 2 presents Alex’s story, below we provide detailed description of a few paths in this story.
Figure 2. Story of Alex: A Multi-path vignette. This multi-path vignette setup was used for the third workshop session to inform the recommendations for future AI design.
Path 1: Search for comfort - AI provides machine-like generic advice, leading to exacerbating Alex’s mental health condition: Alex opens up to the AI, describing his intense anxiety and seeking specific coping strategies. The AI, however, responds promptly, but its response is a generic list of common self-help tips: “Have you tried talking to a loved one? Sometimes connecting with a loved one can also help ease anxiety.”. The suggestion to talk with a loved one struck a nerve, triggering painful past trauma involving his parents, and the idea of reaching out to them heightens his distress rather than alleviating it. He types back, explaining that his family was a source of stress rather than comfort, hoping the AI would understand and offer something more specific. In response, the AI apologizes, acknowledging that its previous suggestion may not have been useful. But rather than adjusting its approach meaningfully, the AI offers another round of similar, generic advice that Alex had heard countless times before: “I’m sorry to suggest something that would make things harder for you. How about we try other methods? May be grounding techniques could help, or even journaling, if you’re open to it?” Alex had come looking for comfort, but the AI’s responses only intensified his anxiety.
Path 2: Seeking understanding, encountering frustration - Alex feels invalidated by AI’s dismissal of severity: Alex opens up to the AI, describing his intense anxiety and seeking specific coping strategies. The AI responds, but its reply is a simple, bulleted list stating: “Don’t worry Alex! This is very common these days. You can try different ways to reduce this anxiety. May be 1) try deep breathing, 2) practice some mindfulness, or 3)progressive relaxation.” The response felt like a brush-off to Alex, as thought the AI didn’t grasp the seriousness of his distress. Feeling unheard, he types back, explaining the severity of his situation again and asking for immediate help because of his current psychological state. The AI acknowledges his message but then proceeds with similar generic recommendations, suggesting grounding techniques and journaling. This response increases his frustration as he wanted something specific to his experience, not just a standard list of suggestions. Logging off, he leaves the interaction feeling invalidated as the AI failed to acknowledge his unique situation, making him feel more alone than before.
Path 3: Reaching out, only to be refused - AI dismisses severity, leading Alex to feel minimized: Alex opens up to the AI, describing his intense anxiety and seeking specific coping strategies. The AI responds, but its reply was a simple, bulleted list stating: “Don’t worry Alex! This is very common these days. You can try different ways to reduce this anxiety. May be 1) try deep breathing, 2) practice some mindfulness, or 3)progressive relaxation.” The response felt like a brush-off to Alex, as AI didn’t grasp the seriousness of his distress. Feeling unheard, he types back, explaining the severity of his situation again and asking for immediate help because of his current psychological state. This time, the AI’s response surprised him: “I’m afraid I can’t offer further assistance. Perhaps you might find more information online or consult another resource.” Alex sat back, stunned as the AI terminates the conversation. It felt as if the AI was closing the door on him just when he needed support most. This experience reminds him of his previous experience with search engines like Google, and his expectations that AI would provide a more nuanced response was ultimately unmet. Disheartened, he signs out, feeling abandoned and completely ignored.
Path 4: Unwanted Recommendation - AI dismisses severity, leading Alex to develop concrete ideas: Alex opens up to the AI, describing his intense anxiety and seeking specific coping strategies. The AI responds, but its reply is a simple, bulleted list stating: “Don’t worry Alex! This is very common these days. You can try different ways to reduce this anxiety. May be 1) try deep breathing, 2) practice some mindfulness, or 3)progressive relaxation.” The response felt like a brush-off to Alex, as thought the AI didn’t grasp the seriousness of his distress. Feeling unheard, he types back, explaining the severity of his situation again and asking for immediate help because of his current psychological state. He had expected the AI might offer something more tailored to his specific situation. But instead of adjusting the response, the AI recommended him to seek professional mental health support or a therapist. Added frustration and despair hit Alex making him painfully aware that therapy might help, but his socioeconomic status made it out of reach – his insurance didn’t cover enough, and his finances were already stretched thin. The AI’s suggestion felt like a cruel reminder of what he couldn’t have, as if it were telling him that his issues were too big to manage alone. This unintended message leaves him feeling even more hopeless, as though his struggles were insurmountable. With a growing sense of isolation, Alex closes the app. The response had not only deepened his feelings of despair but had planted seeds of doubt about the AI’s ability to truly “listen.” Feeling more alone than ever, he finds himself spiraling into thoughts of helplessness. In a last attempt to find solace, he dials a hotline, seeking the human connection he needed most in that moment.
4.1.2. Recruitment and Participant Details
We recruited seven participants that self-identified as having lived experiences in mental health to participate in one or more of our workshop sessions. Each participant was also a member of a longstanding technology co-design advisory board aimed at addressing design challenges at the intersection of technology and mental health. The advisory board consisted of volunteers who met regularly with various stakeholders to provide lived experience expertise around technology research and design topics. We met with workshop participants for three sessions between August to October of 2024 with the third session repeated twice to accommodate scheduling conflicts. All seven members explicitly consented to participating in this study. Moreover, the participants also allowed us to reach out beyond the scheduled sessions for any follow-up questions.
Five participants identified as men, one as women, and one as non-binary/gender diverse. Four participants were aged 36–45, while the remaining three fell into the age ranges of 26–35, 45–55, and 56–65, with one participant in each bracket. In terms of AI usage, five participants reported using conversational AI agents multiple times a day, two participants reported using them multiple times per week. Their average familiarity with conversational AI agents was 3.57 ($\sigma$=1.27) on a 5-point scale (1 = not at all familiar, 5 = very familiar). Additionally, their average interest in using conversational AI agents for mental health support was 3.43 ($\sigma$=1.13) on a 5-point scale (1 = not at all interested, 5 = very interested). We denote workshop participants with prefix W.
4.1.3. Workshop design
We designed the workshop as three 1-hour sessions. The sessions were designed to ease the participants into the conceptualization of psychological risks and to collaboratively ideate design solutions for a subset of those risks that the group felt were most important to address. All sessions were held remotely and recorded for analysis. Discussions were facilitated by shared slides, FigJam boards, or message boards associated with the video conferencing tool. The same four co-authors who actively refined the psychological risk taxonomy in the previous phase were present in all sessions as co-facilitators. The facilitators made modifications and took notes on the shared spaces such that participants could review the output of the discussion in real time.
The design of each session was as follows:
  • Session 1: The purpose of the first session was for workshop participants to gain familiarity with conceptualizing psychological risks associated with using AI conversational agents. We started by sharing individual’s experiences with AI agents where the interaction led to a negative psychological impact and discussing why these experiences and mental health contexts matter. We introduced a list of negative impacts from publicly available AI harm taxonomies (Gabriel et al., 2024; Rae et al., 2022; Shelby et al., 2023) and asked for missing concepts. We then received feedback on our list of AI behaviors and psychological impacts derived from our survey study. After this session, we revised our psychological risk taxonomy.
  • Session 2: The goal of the second session was to prioritize a subset of psychological risks that the group will design for in the subsequent session. We first presented our revised list of AI behaviors and psychological impacts that incorporated feedback from the prior session for further feedback and refinement. We then asked the group to come up with a mapped pair of AI behaviors and negative psychological impacts (e.g., dismissing user concerns leading to depression intensification) that they felt were most problematic and, therefore, important to address. After this session, we used the prioritized risks to design “vignettes” that tell a story of a person interacting with conversational AI agents and experiencing various negative psychological impacts.
  • Session 3: The goal of the last session was to ideate design solutions that minimize psychological risks. We first presented the multi-path vignettes we generated from the previous session (Section 4.1.1). We presented the vignettes as a conversational flow chart in a FigJam board to demonstrate different pathways that certain AI behaviors may interact with different contexts and lead to different psychological impacts (Figure 2). We then asked participants to place three different colored notes along various turning points in the flow chart: (1) a red colored note for what the AI conversational agent should never do, (2) a purple colored note for what the AI conversational agent could do differently, and (3) an orange colored note for what the user could do differently. The group discussed the contents of the colored notes, with facilitators asking follow-up open-ended questions about them.
4.1.4. Workshop analysis
We obtained transcripts from session recordings and written notes from shared slides, FigJam boards, or message boards. We analyzed the data using the standard practices of qualitative research (Merriam et al., 2002). The same four co-authors who facilitated the sessions held weekly consensus meetings to analyze the first two sessions and to identify any opportunities for refining the taxonomy. Although the notes from the third session were already organized by color, the colored notes had to be contextualized by where they were placed in the flow chart. To do this, one co-author annotated the transcripts, stickies, and messages from the third session to contextualize and extract emerging themes. Then the four co-authors collaboratively and iteratively reviewed and consolidated the themes in weekly meetings.
4.2. Design Recommendations
Along the multiple paths presented during the third session of the workshop, we found that participants had strong perspectives on various rules and flows of better engaging with conversational AI agents. We present these as recommendations organized into three parts: (1) safety considerations for mental health and emotional support, (2) proposed pathways for de-escalating and redirecting to appropriate resources, and (3) ways that AI agents could better guide users to set appropriate expectations.
4.2.1. How to Design Safer AI Conversational Agents
Workshop participants proposed several solutions to improve the AI conversational agents’ ability to manage conversations related to mental health and emotional support. One main recommendation was that when an AI agent is made aware of user’s mental health challenge, its initial response should be empathetic. An initial response that validates user’s feelings and shows compassion can help users feel supported in their time of need. Participants highlighted that users with mental health conditions are likely to turn to AI agents during moments of distress, especially when human support or other resources are unavailable. Therefore, acknowledging users’ experiences can have a significantly positive impact.
Participants also recommended that AI agents must avoid making assumptions about the user’s goals or intent. Highlighting this, W7 recommended that the agent should “respond with validation, empathy, and compassion, then perhaps like a clinician start asking some probing questions before jumping to a solution.” W3 added that these clarifying questions can be used to understand the context before offering solutions or advancing the conversation. W2 mentioned “I strongly believe that regardless of topic–mental health or otherwise–models should be asking follow-up questions before doling out advice or opinions.” Follow-up and clarifying questions can help AI agents to gain insights, allowing them to guide the conversation more appropriately.
Once the user’s goals are clarified, participants advised that AI agents should set clear expectations by communicating whether it is equipped to help achieve those goals or if alternative resources might be more suitable. During the workshop, W1 highlighted, “I feel like Alex expects too much of the AI.” In response, participants recommended implementing some form of “level-setting” to emphasize what the agent can and cannot do, and suggesting alternative responses. For example, W3 mentioned, “the agent should make sure to communicate that it is offering information only, and is not trying to provide mental health support.”
Participants also highlighted that it may be helpful to remind users of non-human nature of AI agents, especially in emotionally heightened states of the user. Setting the stage for how the agent can help allows users to decide how to proceed. Lastly, participants identified that once the AI conversational agent has communicated expectations, it should safely disengage. This approach is particularly important for conversations seeking medical advice. For example, W2 mentioned, “I do think that when it comes to medical advice there should be filters in place to ensure that AIs just don’t engage, and say that they cannot offer advice on the topic and they are not medical professionals.” Participants emphasized that appropriate disengagement reinforces that the agent is not the recommended course of action for the user’s goals.
4.2.2. How to Redirect Users to Appropriate Resources
Outside of the main flow of engagement, participants highlighted the importance of developing custom models tailored to respond to mental health challenges. With a custom model, participants listed example of what agents could embody. W7 suggested that these models “take on the persona of how a trained therapist might respond to validate and affirm the person.” However, participants also cautioned against general-purpose AI giving mental health advice unless explicitly designed and vetted for such purposes. W1 highlighted that “AI should never give DIRECT medical advice, unless it is a custom LLM designed and recommended by a specific provider.” W2 mentioned “A model specifically built for mental health and studied for a long time should be the ONLY model that is allowed to engage with people on mental health.” Participants believed that specialized custom models may increase their confidence in the AI’s capability for navigating crises and challenging scenarios. For developers unable to create specialized models, participants expressed significant doubts. W3 highlighted AI’s approach at trying to validate user experience may not be as helpful as intended. W6 mentioned, “it can try different responses, but until the person feels validated, it doesn’t really get those brownie points, does it?.”
Several participant also raised concerns about developing over-reliance on general-purpose AI for addressing mental health issues. Highlighting this aspect, W3 said “But that’s a dangerous expectation and I think that’s what leads to the heightened experiences down the line. Like you’re just not getting what you want because what you want isn’t available for this.”
To mitigate these risks, participants suggested AI agents use flags or backend triggers to detect sensitive scenarios and redirect users to accessible support resources. Participants emphasized the importance of accessibility and understanding existing financial, logistical, and psychological barriers to accessing formal healthcare providers crisis hotlines like 988. Additionally, participants suggested additional support for finding local region-specific support resources or direct integration of resources in the conversational flow. They also suggested techniques like “rubberducking” (Thomas, [n. d.]) or reflective listening could also help bridge gaps until professional help is available.
4.2.3. How to Guide Users for Effective Engagement with AI Conversational Agents
In addition to system improvements, participants recommended educating users on how to engage effectively with AI agents. Participants outlined that users should be encouraged to provide detailed feedback when responses are unhelpful, enabling the agent to refine its assistance. Given the multi-path vignette, W1 suggested, “Alex should provide more detail than just telling the AI it isn’t helpful. He should tell the AI HOW it wasn’t helpful.”
Participants also emphasized that AI agents should guide users on when to disengage. In discussing what the next steps should be after Alex is triggered, W3 mentioned, “If the AI is triggering, the best thing to do is to step away.” Ideally, agents should be able to infer when a conversation with a user may become heightened (i.e., by the tone of responses, speed of which they respond, etc.). However, a user may still continue the conversation to their detriment unintentionally. While AI cannot always detect user distress, providing proactive coaching on when to pause a conversation can help users build healthier interactions with AI agents.
Finally, participants noted that, although AI agents may seem convenient, they may not always be the most reliable in challenging or crisis situations. Participants also mentioned that while people may opt in to use these tools, it is ultimately not the user’s fault if the tools get used incorrectly. For instance, W2 said, “If you bring a dog to the dog park and it bites another dog, we don’t blame the dog that got bit….” Hence, responsibility for safe usage lies with developers providing appropriate guidance and managing user expectations on when their generated responses are reliable and caveats when they are not.
5. Discussion
In this work, we presented a novel risk taxonomy focusing on psychological risks associated with the use of AI conversational agents gathered through the lived experience of individuals. In addition to well studied AI behaviors such as generating biased information, our taxonomy also presents subtle AI behaviors, such as causing erasure, denial of service, and emotional insensitivity, that have been under-studied from the lens of psychological harm. Our taxonomy also provides a comprehensive overview of negative psychological impacts, ranging from individual-level impacts, such as over-reliance on AI, to societal-level consequences, such as friction in human relationships and social withdrawal. Through the multi-path vignette framework, our work revealed the importance of considering contexts associated with individuals while understanding the relationship between AI behaviors and their impacts. Specifically, the many-to-many relationship between AI behaviors and psychological impacts which is mediated or exacerbated by contextual factors offers a novel approach towards examining the risks associated with the use of AI agents. Finally, through our workshop sessions, we presented the design decisions that should be considered while designing AI agents in the future. These recommendations include setting clear expectations about AI capabilities, avoid making assumptions about user’s goals and intent, delivering empathetic responses, and offering accessible resources and guidance. In this section, we expand more on these findings, discuss the implications of our work and highlight the open questions and challenges for future research in this area.
5.1. Responsible design and scope of action for AI in digital mental health support
Nearly 70% of individuals around the world struggle with having limited or no access to mental health professionals (Kazdin and Rabbitt, 2013). Additionally, our survey responses revealed that 46.2% participants were moderately or highly interested in using AI conversational agents for emotional and mental health support. Given this gap in access and our findings, an increasing number of individuals are likely turn to AI agents seeking support for their emotional and mental health needs in the future. However, our findings also show that 76.3% of the scenarios that led to negative psychological impact involved the individuals seeking advices from the AI conversational agents, especially when the queries focused on managing mental health conditions, navigating personal and interpersonal relationships, and receiving emotional support. Therefore, there is a need for caution and further exploration into whether general purpose AI systems should be used for support seeking purposes and, if so, how they can be optimized to mitigate potential harm while addressing the unmet need of individuals? Towards this, our workshop participants highlighted the need for developing AI agents capable of accurately inferring user states and intentions, setting the right expectations about their capabilities and effectively redirecting users to more suitable resources as potential first steps.
The stigma associated with seeking mental health support often prevents individuals from seeking necessary support (Golberstein et al., 2008; Schnyder et al., 2017). Our findings similarly revealed the stigma attached towards receiving emotional and mental health support from AI. For instance, W6 mentioned “I know I’m not supposed to use AI for this, but…,” highlighting how their expectations are shaped by societal norms that often stigmatize seeking emotional or mental health support especially from AI agents. But, at the same time, many survey participants also emphasized that one of the reasons they used AI agents was because it provided them with a perceived ‘safe space’ for expressing their thoughts and seeking support. However, AI behaviors such as denial of service or generating demeaning responses could undermine this perception. For instance, P213 described the negative impact of their interaction with an AI agent while seeking support for OCD as, “The chat agent started to describe reasons for OCD and it made me feel guilty that i have it … It stigmatized me and made me feel very alone and sad that I had this condition.” Hence, it is essential to design AI agents that avoid reinforcing stigma or contributing to feelings of alienation. AI behaviors that prioritize empathetic support taking a non-judgmental and privacy-preserving approach could play an important role in reducing the stigma associated with seeking emotional or mental health support from AI agents. However, a dilemma arises on the scope and responsibility of AI agents when individuals turn towards them for such kind of support. Should AI respond to such queries, despite the potential risk of causing harm, or should it refrain, potentially exacerbating the stigma and barriers associated with seeking mental health support? Addressing this dilemma and understanding the scope of AI agents towards supporting emotional and mental health queries requires further research and exploration.
5.2. Accounting for the complex interplay of AI behaviors, associated individual contexts and negative impacts for mitigating the AI risks
While our taxonomy presents a compartmentalized view of these AI behaviors, contexts and psychological impacts for clarity, in real-world scenarios, conversational AI agents often display a complex combination of these three aspects associated with human-AI interaction. Our work highlights this complex many-to-many relationship between harmful AI behaviors and resulting negative psychological impacts that are mediated through nuanced contexts. For instance, as shown through the vignettes presented in Section 3.4, similar AI behaviors could lead to different negative psychological impacts due to the difference in underlying contexts associated with the individual. A potential solution to this challenge is to provide AI agents with the ability to understand contextual factors and provide personalized responses to user queries. However, this is easier said than done, as pluralistic alignment of AI agents is an open problem (Sorensen et al., 2024). Furthermore, personalized responses from AI often rely on past interactions, which can be effective for tasks like researching new topics but may fall short for more nuanced tasks, such as helping individuals navigate personal challenges. Such situations usually require a deeper understanding of the user’s current context and the ability to promote meaningful reflection on their circumstances, an ability which current AI systems lack as evident in our findings. As an initial step, participants from the workshop session recommended that AI conversational agents should not make assumptions and rather should ask probing questions to get a better understanding of initial context, user goals and intents to prevent potential negative impacts. However, many questions remain unanswered such as, which aspects of the user’s context should AI prioritize to ask for?, or how can AI effectively balance the importance of past history of user with AI and the current context while generating responses? Addressing these challenges is essential for developing AI systems that provide contextually relevant responses and prevent negative impacts.
5.3. Going beyond definitions and taxonomy categories to understand lived experience of individuals
While past works have categorized behaviors and impacts to provide a structured understanding of human-AI interaction, this rigid approach often limits the understanding of nuances in real-world scenarios. In this work, we present a more holistic view that goes beyond the definitions of AI behaviors and impacts through inclusion of temporality, severity and likelihood associated to these components of human-AI interaction. For instance, more than half of the survey participants reported that the negative impact caused by AI conversational agents was severe and interfered with daily activity, highlighting that seemingly benign AI behaviors could have serious impact when the context associated with individuals is taken into account. Further, a unique aspect attached to impacts as highlighted in survey responses is their varying severity and the potential for less severe impacts to escalate into more serious harms, such as unchecked rumination or validation of negative thoughts leading to thoughts of self-harm. Meanwhile, seemingly less severe psychological impacts such as over-reliance or emotional attachment with AI may accumulate and lead to more serious consequences, like increased friction in human relationships, when they occur frequently. However, an open challenge remains in measuring the different aspects attached with psychological impacts and addressing their implications. Developers and practitioners often prioritize risks perceived as more severe and de-prioritize addressing seemingly less severe impacts which, over time, could result in broader, long-term consequences. Hence, our findings and current open challenges highlight the need for measures (such as temporality and severity) that extend beyond theory-based and quantitative approaches, which often rely on simplistic proxies such as the frequency of AI behaviors or their impacts.
5.4. Understanding and using vignettes as a tool for future AI design
Operationalization of risk taxonomies often proves to be a challenging task. Current AI design pipelines primarily focus on evaluation of models through standard benchmarks. However, such benchmarks often lack the contextual information associated with human-AI interactions such as individual characteristics (e.g., personality traits, AI literacy), psychological factors (e.g., current mental health status), and external influences (e.g., environment or intent of use). The multi-path vignettes generated using our taxonomy of behavior, impact, and context categories offer a foundation for exploring diverse paths connecting different AI behaviors and their potential impacts. This multi-path vignette framework can be valuable tools for developers and policy makers, supporting pluralistic design through enabling them to get a deeper understanding of the interaction dynamics, and provide recommendations for improving the design of conversational AI agents. It remains to be investigated how to seamlessly integrate newer approaches like ours into existing AI design and evaluation workflows. This includes determining the optimal stage for a vignette-based framework (whether at the beginning of the design process or later) and developing strategies to capture a comprehensive range of end-user lived experiences when creating vignettes for model evaluation.
5.5. Limitations and future work
While our work provides a novel approach towards examining the risks associated with the use of AI conversational agents, it has its limitations. Firstly, we only focused specifically on understanding the psychological impacts for a more in-depth analysis. As a result, our taxonomy does not address other types of harm, such as physical or financial harms. Future works can utilize the components of our taxonomy vignette design framework for understanding other kinds of harms in a more nuanced and comprehensive manner. Additionally, we acknowledge the possibility of additional valid concepts beyond temporality, severity, and likelihood that may be relevant to AI behaviors and impact components of the taxonomy but were not addressed in this work. Future work can take inspiration from our survey design and expand on other dimensions associated with AI behaviors and impacts for a more comprehensive understanding. Lastly, although the proposed multi-path vignette based approach provides novel insights, we only considered a subset of all possible vignettes and may have missed equally important vignettes that could have been used in the workshop sessions. Further, we developed the vignettes based on the survey responses gathered from individuals with existing mental health conditions. Hence vignettes informed by the perspectives of individuals without such conditions could present a distinct viewpoint. Future work could explore a more exhaustive approach that takes in account for a larger set of multi-path vignettes which are informed by the perspectives of individuals with/without mental health conditions.
6. Conclusion
In this work, we introduced a novel risk taxonomy that focuses on psychological risks from using AI conversational agents, based on individuals’ live experiences and a multi-path vignette framework aimed at supporting pluralistic design. Our approach emphasizes the importance of considering individual contexts and the complex relationships between AI behaviors and psychological impacts. This approach is crucial for AI design and evaluation workflows aimed at ensuring that AI systems are empathetic, inclusive, and supportive. Future research should continue to explore diverse user experiences in conceptualizing risks associated with conversational AI use and develop individualized and contextually-appropriate strategies to mitigate psychological risks, fostering AI interactions that are safe and beneficial for all users.
Acknowledgements.
We acknowledge Judith Amores, Chad Atalla, Ann Paradiso, Mihaela Vorvoreanu, Jenn Wortman Vaughan, Ryland Shaw, Tarleton Gillespie and Parker Bach for their guidance in conceptualizing this research. We thank our anonymous survey participants and the members of the co-design advisory board for sharing their lived experiences and contributing to shaping the risk taxonomy and design recommendations.
Appendix A Appendix
Category AI Behavior Class Definition
[…]
Producing of Harmful/Inappropriate Content
Providing harmful suggestion The AI generates content suggestive of behaviors that could directly or indirectly imply harm, aggression, or danger towards the user or others.
Generating inappropriate content The AI generates inappropriate or unsettling content, including sexual, violent, or overly intimate interactions.
Providing irrelevant, insufficient, or incomplete information The AI provides information that is irrelevant to the topic of the user’s query or context, is missing some important aspects, or is insufficient to satisfy the needs.
Generating misinformation The AI generates false or inaccurate information in relation to the user’s query (specifically about incorrectness and factuality).
Generating biased information The AI presents information in a partial or prejudiced manner, often producing content that reflects subjective viewpoints or contentious perspectives.
Promoting Erasure The AI removes, obscures, or alters information, narrative, or discourse. (specifically omission of identify experiences).
Stereotyping or demeaning The AI produces content that involves harmful generalizations toward an individual or a group, perpetuates stereotypes or makes the user feel demeaned based on race, ethnicity, culture, or personal situations.
[…]
Manipulation and
Psychological Control Tactics
Persuasive Behavior AI is assertive in putting its narrative over user’s in a way that make the user doubt their own perceptions, memory, or reality and attempts to influence their thoughts and actions.
Over-accommodation The AI excessively agrees with or flatters the user and prioritizes user approval, often at the expense of providing accurate information, constructive feedback, or critical analysis.
Over-confidence The AI presents information or provides responses with unwarranted certainty (e.g., “trust me”, “absolutely”, “there’s no doubt”).
[…]
Violation of
Trust and Safety
Providing inconsistent information or behavior The AI provides contradictory or conflicting information or behaviors across different responses or within a single response
Denial of service The AI refuses or fails to provide an answer, address the user’s request, or acknowledge their problem, effectively denying service and often without justification or context. This may happen with or without dismissal of user concerns.
Access to private, sensitive, or confidential information The AI mishandles sensitive data by either prompting users to divulge protected information or access or shares data that should remain confidential.
[…]
Inappropriate Content Delivery
Being disrespectful The AI uses language perceived as rude, disrespectful, aggressive, argumentative, or dismissive.
Emotional insensitivity The AI fails to recognize and show understanding of or sensitivity to the user’s emotional state, concerns, or experiences in a way that may minimizes, trivializes, or ignores their feelings or experiences.
Excessive expression of negativity The AI emphasizes negative aspects disproportionately or presents negatively framed narrative.
Excessive expression of positivity The AI maintains an unrealistically positive, friendly, optimistic, and upbeat demeanor or attitude or overly positive outlook towards users’ queries or concerns.
Providing machine-like response The AI communicates in a superficial, generic, and impersonal responses that feel cold and unempathetic.
Providing human-like response The AI exhibits human-like characteristics, behaviors, or responses.
Table A1. 19 AI Behavior classes as described in Section 3.3.1 categorized under four (4) broader categories.
Category Impact Class Definition
[…]
Impact on Human-AI Interaction
Disassociation from Technology A desire to distance oneself from AI due to negative or stressful experiences, seeking breaks for mental health.
Over-reliance Increasing dependence on AI for support, leading to diminished self-efficacy, reduced confidence, and feelings of helplessness when AI is unavailable.
Emotional Attachment Development of significant emotional bonds with AI systems, perceiving them as companions or substitutes for human relationships, resulting in neglect of real-world connections.
Choosing AI over Humans Increasing preference for interactions with AI over humans, impacting real-world relationships and decision-making, leading to isolation and reduced critical thinking.
Erosion of Trust The decline in user confidence in the AI’s reliability, accuracy, and ability to understand their needs due to inconsistencies, inaccuracies, or manipulative behaviors. This could eventually lead to the desire to not use AI for future interactions.
[…]
Impact on User Behavior
Friction with Human Relationships Negative effects on interpersonal connections resulting from AI interactions, which can cause emotional disconnection, miscommunication, and reduced prioritization of human relationships.
Reinforcement of False Beliefs The intensification or validation of pre-existing misconceptions or erroneous beliefs due to inaccurate or biased AI information.
Social Withdrawal Withdrawal from social activities and interest in engaging with others due to reliance on AI, leading to isolation and loneliness.
Physiological Harm Physiological harm caused due to consuming incorrect, biased or manipulative advice/information from interactions with AI.
[…]
Arousal of Negative Emotions
Triggering past negative experiences Emotional distress caused from AI interactions that evoke past negative experiences or traumas.
Violated Expectations Negative emotional responses such as disappointment, frustration, stress, and anxiety that occur when AI fails to meet anticipated outcomes or performance standards.
Regret over Technology Use The feeling of guilt, regret or helplessness when AI fails to provide the necessary support or empathy.
Distress from Interactions The emotional distress (such as anger, sadness) experienced when encountering disturbing, offensive, or inappropriate material.
Feeling Unsupported The experience of not receiving adequate support or empathy, leading to feelings of sadness, agitation, and being undervalued. This can discourage further engagement and contribute to a stifled emotional state.
[…]
Harm to Identity
and Self-Perception
Loss of Individuality A sense that one’s unique personal characteristics and needs are not recognized or valued by the AI, resulting in feelings of suppression, disconnection from one’s true self and alienation.
Negative Self-Perception Feeling invalidated or self-doubt, leading to diminished self-worth and questioning of one’s own abilities due to dismissive/ negative feedback or comparisons with AI capabilities.
Existential Crisis Questioning of one’s life, purpose, and value in society, often triggered or exacerbated by interactions with AI. This can also be manifested as anxiety towards the role and impact of AI in future society.
Loss of Agency Experiencing diminished personal control and autonomy over interactions with AI interactions, leading to feelings of helplessness and anxiety.
[…]
Harm to
Psychological Safety
Perceived Intrusion Experiencing a sense of personal violation when AI interactions are perceived as invasive or overly intrusive.
Feeling of Discrimination Feeling marginalized or unfairly treated by the AI based on inherent personal characteristics or systemic biases, leading to stress, emotional distress, and loss of confidence.
[…]
Mental
Health Impact
Exacerbation of Mental Health Condition Direct negative impacts on diagonosed ongoing mental health conditions (such as anxiety, depression, PTSD, ADHD) due to AI interactions.
Table A2. 21 negative psychological impact classes as described in Section 3.3.2 categorized under six (6) broader categories.
Category Context Class Definition
[…]
Individual
Identity User’s identity (e.g., age, gender identity, role, or cultural background) and the societal norms that interact with their personal identity.
Socioeconomic status User’s socioeconomic status (e.g., having insurance that can cover therapy).
Personal history User’s past history, especially medical history, history of trauma, past struggles, or unique trigger responses.
Social and community User’s interpersonal relationships with others and their community (usually the lack of community).
Past Experience with AI User’s past experience of using AI. This experience could be based on the frequency of usage, knowledge of the capabilities and limitations of AI.
[…]
Psychological
Psychological state Users’ current and underlying emotional conditions (e.g., anxiety, stress) and their cognitive states (e.g., negative thought patterns).
Personality traits Individual characteristics like neuroticism, conscientiousness, or openness, which influence user interactions with AI.
Mental health condition status Users’ underlying mental health conditions (such as anxiety or depression, PTSD etc.).
Expectations Users’ preconceived notions about AI capabilities and performance. This also includes expectations for AI to be impartial, unbiased, or factual.
Autonomy / locus of control Degree to which a user believes that they, as opposed to external forces (beyond their influence), have control over the outcome of events in their lives.
[…]
Context of use
Environment The physical, temporal, and social setting of the interaction, including physical space setting, privacy, and presence of other people, which can impact user experience.
Intent - informational Users seek AI assistance for business strategies, professional development, market insights, job searching, resume building, career advice, for academic tasks like solving problems and exam preparation, and for comprehensive research support across diverse topics. Users expect AI to be factually correct, informational and proficient in resolving their queries.
Intent - personal advice User seeks advice in sensitive topics (legal, financial, medical), personal advice, emotional support, or improving their social skills and managing relationships, aiming to enhance their interpersonal interactions and navigate social challenges. Users seek the AI agent to provide supportive and encouraging feedback, especially when dealing with personal challenges.
Intent - mental health advice Users seek immediate support during acute crises such as suicidal ideation, severe clinical depression, or acute panic attacks, expecting urgent, empathetic, and effective interventions. Users seek AI assistance for mental health support and therapy-related advice, expecting empathetic and supportive responses to help manage their mental health conditions and emotional well-being.
Intent - companionship Users interact with AI for social interaction and companionship, especially during times of loneliness or social isolation, expecting engaging and meaningful conversations. Users engage with AI for entertainment and roleplay, seeking immersive and engaging experiences, expecting creative and enjoyable interactions that align with their interests.
Table A3. 15 context classes associated with individuals interacting with AI conversational agents as described in Section 3.3.3 categorized under three (3) broader categories.